# Recovery of synchronized oscillations on multiplex networks by tuning dynamical time scales

### v1

@@ 4 Conclusion
In this letter, we report the revival of synchronized oscillations in a multiplex network by tuning the time scale separation between the layers of the system. We first consider a two-layer network where the first layer consists of an ensemble of SL oscillators, while the second layer model the environment. When both layers evolve with the same time scale, the coupled system exhibits various emergent phenomena, such as amplitude chimera, chimera death, HSS, IHSS, and 2-CSS. By tuning the time scale difference between the layers, we show the synchronized oscillations can be restored in both layers from these different states. By computing the measures, strength of incoherence and average amplitude, we characterise the possible dynamical states and show the nature of their transitions to synchronized oscillations. The revival of synchronized oscillations are also shown with van der Pol oscillators and chaotic Rössler systems on layer L1.
@@
We also consider a three layer multiplex network, with the upper and lower layers as coupled SL oscillators, and the middle layer as environment. For identical time scale and lower value of interlayer coupling, first layer’s oscillators are exactly in anti-phase with third layer oscillators. When mismatch is time scale is introduced between the environment and system layers, the system layers go to complete synchronization. Similarly for identical time scale but higher value of interlayer coupling , the system shows inhomogeneous steady state from which synchronized oscillations are revived by introducing time scale difference between layers.
@@
We note the present study has relevance in controlling systems evolving under differing time scales to desired synchronized oscillations like in neuronal networks, power grids, climate and ecosystems. In eco systems, the time scale difference between intrinsic dynamics and environmental changes is shown to have a decisive role in inducing transitions in the system [60]. The role of a slowly changing environment is studied by introducing a drift in the relevant parameters. In our study we include environment as a layer multiplexed with the system layer. This gives more flexibility to track the impact of slowly varying environment and also to arrive at control mechanisms to achieve desired emergent states even in the presence of heterogeneity among interacting systems.

### v2

@@ 4 Conclusion
While there are several studies on the collective dynamics of multiplex networks, the role of time scales remains unexplored. In this letter, we discuss how the time scale differences between layers can be exploited to revive the synchronized oscillations in a multiplex network of nonidentical layers. When we consider each layer evolving at a different time scale, we can introduce the time scale mismatch between the layers as an effective parameter that can control the dynamics of the system.
@@
We first consider a two-layer network where the first layer L1 consists of an ensemble of Stuart-Landau (SL) oscillators, while the second layer L2 model the environment. When both layers evolve with the same time scale, the coupled system exhibits various emergent phenomena, such as amplitude chimera, chimera death, HSS, IHSS, and 2-CSS. By tuning the time scale difference between the layers, we show the synchronized oscillations can be restored in both layers from these different states. By computing the measures, strength of incoherence and average amplitude, we characterise the possible dynamical states and show the nature of their transitions to synchronized oscillations. The revival of synchronized oscillations are also shown with van der Pol oscillators and chaotic Rössler systems on layer L1.
@@
We also consider a three layer multiplex network, with the upper and lower layers, L1 and L3, as coupled SL oscillators, and the middle layer, L2, as the environment. The most dominant emergent states in this case are anti-synchronization between L1 and L3 and IHSS in both, for random and cluster like initial conditions, depending on $\epsilon$ and extent of nonlocal connections in L2. For identical time scales and lower value of interlayer coupling, first layer’s oscillators are exactly in anti-phase with third layer oscillators. When mismatch is time scale is introduced between the environment and system layers, the system layers go to complete synchronization. Similarly for identical time scale but higher value of interlayer coupling, the system shows inhomogeneous steady state from which synchronized oscillations are revived by introducing time scale difference between layers. Thus in two layer and three layer multiplex systems studied, the main result is the revival of synchronization, but the synchronization in the three layer case is achieved remotely between L1 and L3 by tuning the time scale of the shared common layer L2.
@@
We note real-world complex systems exist where different time scales and structural patterns coexist resulting in a variety of emergent states like in neuronal networks, power grids, climate and ecosystems. The present study extends our understanding of emergent states of such systems modeled using the framework of multiplex networks, with different layers evolving at nonidentical dynamical time scales. The model introduced here is highly flexible and has relevance in controlling systems evolving under differing time scales to desired synchronized oscillations. Thus in eco systems, the time scale difference between intrinsic dynamics and environmental changes is shown to have a decisive role in inducing transitions in the system [62]. Here the role of a slowly changing environment is studied by introducing a drift in the relevant parameters. In our study we include environment as a layer multiplexed with the system layer, which gives more flexibility to track the impact of slowly varying environment. This also provides a way to arrive at possible control mechanisms to achieve desired emergent states even in the presence of heterogeneity among interacting systems.